I hope this question is under the right category. The lesson is called conics.
Explain how the difference in the equations makes the two graphs different. Same question for all problems listed. Thank You
A. (Y+3)^2/23 - (X+2)^2/17=1
(x+2)^2/17 - (y+3)^2/23=1


B. (y-4)^2/7 + (x+9)^2/3=1
(x-4)^2/7 + (y+9^2/3=1


C. (Y+7)^2/11 - (X-4)^2/15=1
(Y+7)^2/15 -(X-4)^2/11=1

Answer by Alan3354(69443)

I hope this question is under the right category. The lesson is called conics.
Explain how the difference in the equations makes the two graphs different. Same question for all problems listed. Thank You
A. (Y+3)^2/23 - (X+2)^2/17=1
(x+2)^2/17 - (y+3)^2/23=1

B. (y-4)^2/7 + (x+9)^2/3=1
(x-4)^2/7 + (y+9^2/3=1

C. (Y+7)^2/11 - (X-4)^2/15=1
(Y+7)^2/15 -(X-4)^2/11=1
----------------------
DL the (free) software at www.padowan.dk.com/graph/
Graph these, check them out.
Use Insert, Insert Relation